OpenStudy (anonymous):

Solve sqrt3 tanx - 1 = 0

OpenStudy (anonymous):

I thought we solved this question 2 days ago... Solving it again? :)

OpenStudy (anonymous):

Here's a video for your exact question (It's 45 seconds)! Source: http://www.youtube.com/watch?v=JShnLEvb8Ac

OpenStudy (anonymous):

\[\sqrt{3}\tan x -1=0\] We are trying to isolate x so our first optiion would be to move the one over \[\sqrt{3}\tan x =1\] Then we would divide both sides by square root 3 \[\tan x = \frac{1}{\sqrt{3}}\] Inverse both sides ... \[\ x =\tan ^{-1} (\frac{1}{\sqrt{3}})\]

OpenStudy (anonymous):

\[\huge \sqrt 3 \tan x-1 =0 \rightarrow \sqrt 3 \tan x =1\] \[\huge \rightarrow \tan x =\frac{1}{\sqrt 3}\] \[\huge \rightarrow \tan x =\tan 30^0 \rightarrow x= 30^0\] \[\Huge \tan 30^0 = \frac{1}{\sqrt 3}\]